### v2

# LaDTalk: Latent Denoising for Synthesizing Talking Head Videos with High Frequency Details

Abstract
Identity-specific audio-driven talking head generation (THG) is crucial for applications in filmmaking and virtual reality. Despite significant progress in existing end-to-end methods, they often struggle to generate videos with high-frequency details due to limited expressivity. In this paper, we tackle this challenge by introducing LaDTalk, a framework designed to generate photorealistic talking head videos. Specifically, we redefine the task of ID-specific THG as temporally consistent face deblurring, aiming to restore audio-synchronized but blurred faces from one-shot base models. This approach allows LaDTalk to leverage the strengths of both one-shot methods and ID-specific approaches, enabling the synthesis of lip-synchronized talking head videos with high fidelity and detailed textures. To achieve this, we theoretically establish the noise tolerance of Vector Quantised Auto Encoders (VQAEs) using Lipschitz Continuity theory. Building on this analysis, we introduce a plug-and-play Space-Optimized VQAE (SOVQAE) that recovers fine-grained textures while ensuring temporal consistency. Extensive experiments demonstrate that LaDTalk achieves state-of-the-art performance in video quality and out-of-domain lip synchronization accuracy when integrated with Wav2Lip as the base model. Furthermore, SOVQAE enhances the performance of various base models, confirming its versatility and effectiveness.
1Beihang University  2Psyche AI Inc.  3The University of Alabama at Birmingham  4CASIA   5Tencent
1 Introduction
The generation of photo-realistic, speech-driven ID-specific talking head videos has significant potential across various domains, including filmmaking [41], virtual reality [42], and digital avatar creation [43]. The goal of this work is to synthesize high-fidelity talking head videos that achieve precise lip synchronization while preserving fine-grained details such as hair, facial wrinkles, moles, eyelashes, and the intricate contours of the lips. These high-frequency details are critical for enhancing realism in synthesized videos.
One-shot techniques [1, 12, 11, 44, 51] have made significant progress in lip synchronization but often struggle to maintain temporal consistency and visual fidelity, leading to issues such as tooth flickering and inconsistent lip thickness. These inconsistencies arise from multi-ID training paradigms that rely heavily on large-scale datasets [10, 56].
To address these challenges, NeRF-based approaches [2, 3, 39, 14, 52] have been introduced, demonstrating impressive results in preserving high-fidelity identities. However, these methods, which are primarily based on Multilayer Perceptrons (MLPs), encounter difficulties in capturing high-frequency details due to the phenomenon known as “spectral bias” [5, 6, 59]. Recent studies on the neural tangent kernel (NTK) [57, 4] reveal that spectral bias arises from the pathological distribution of NTK’s eigenvalues, where most eigenvalues are very small, limiting the convergence speed for high-frequency components.
In contrast, our work focuses on addressing the challenge of generating high-frequency details in ID-specific talking head videos. Specifically, we introduce LaDTalk, an effective framework designed to enhance the output quality of Wav2Lip [1]. By treating its outputs as blurred approximations of ideal results, we redefine the task as one of temporally consistent face deblurring. This design leverages the robust audio-lip synchronization capabilities of one-shot models while incorporating advanced restoration techniques for high-frequency details.
A key challenge in this approach lies in ensuring temporal consistency during denoising while preserving high-fidelity details. To address this challenge, we analyze the forward dynamic process of Vector Quantised Auto Encoders (VQAEs) [^1] We use the term VQAE to distinguish it from VQGAN [21], which combines a VQAE with a transformer network and emphasizes image synthesis. Here, we focus on the auto-encoder nature of VQAE. in the latent space using the Lipschitz Continuity theory of neural networks [9, 8]. Based on this analysis, we theoretically prove the noise robustness of VQAEs (see Sec. 3). This novel theory demonstrates that VQAEs can preserve high-frequency detail information within a discrete codebook while exhibiting noise tolerance, as illustrated in Fig. 2. Building on this insight, we propose a codebook regularization loss to enhance the noise robustness of VQAEs (termed Space-Optimized VQAE, SOVQAE) and facilitate temporally consistent face deblurring.
Figure 1: (a) Illustrates Voronoi regions in 2D and explains the denoising mechanism in VQ space. Given a mapping point $P$, there exists an open disk $\{{\bf x}\in V_{A}\cup V_{B}\cup V_{C}:\|{\bf x}-P\|_{2}<s\}$, where all points in the disk share the same anchor point $A$. (b) Demonstrates the robustness of VQAE under slight input disturbances. When image perturbations occur, the VQ mechanism correctly matches the appropriate codebook vector in the latent space, provided the affected vector remains within the correct Voronoi region (e.g., the yellow-red vector in the figure).
More importantly, SOVQAE is a plug-and-play model that can be integrated into various base models to enhance their performance. Our experiments show that integrating SOVQAE improves the visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Especially, when integrated with Wav2Lip [1], LaDTalk surpasses state-of-the-art (SOTA) methods in terms of lip synchronization accuracy, video clarity, and high-frequency detail preservation. Comprehensive evaluations highlight LaDTalk’s superiority in multiple metrics, making it a promising solution for high-fidelity talking head generation.
In summary, this work makes the following contributions:
  • To our knowledge, this study pioneers the exploration of noise robustness in VQAE through the lens of Lipschitz Continuity theory.
  • We propose SOVQAE, a simple and effective plug-and-play model for ID-specific talking head generation that enhances high-fidelity and high-frequency details, supported by rigorous theoretical analysis.
  • Through extensive experiments, we demonstrate the validity and superiority of LaDTalk over SOTA methods across multiple metrics.
2 Related Work
2.1 Audio-Driven Talking Heads Generation
One-shot Methods, as delineated in [1, 11, 12, 44, 51, 15, 53, 54], harness a plethora of talking video clips to capture the generalizability across multiple faces, thereby attaining commendable lip-synchronization performance. For instance, Wav2Lip [1] pioneers the integration of a lip-sync expert to synchronize audio segments with generated lip movements. DINet [12], in accordance with audio features, induces spatial deformations on reference image feature maps, facilitating few-shot generation capabilities. Despite these advancements, these methods continue to grapple with maintaining identity consistency across video frames. To counter this limitation, ID-specific NeRF-based methods [2, 39, 14, 3, 52, 55] are proposed, focusing exclusively on identity-specific data for training. Although this approach mitigates issues of identity instability, the constrained volume of training data hampers their cross-audio synchronization capabilities. To this end, GeneFace [14] and SyncTalk [2] enhance their audio generalization performance by leveraging pre-trained audio-to-landmarks predictors and audio encoders [1] respectively. Although these initiatives successfully bolster the cross-audio generalization of NeRF methods, instances of cross-audio desynchronization persist. Building upon the foundational capabilities of Wav2Lip within our pipeline, LaDTalk achieves superior cross-audio synchronization performance characteristic of one-to-one methods. Moreover, our approach also excels in visual quality, attributable to the temporally consistent latent denoising prowess of our SOVQAE.
2.2 Face Restoration
Blind Face Restoration [30, 20, 31, 32, 33, 34, 35, 36, 37] endeavors to rectify diverse degradations in facial imagery captured in non-ideal conditions, including compression, blur, and noise. Given the unknown nature of the specific degradations, this task is inherently ill-posed. The prevalent approach to address this challenge is to integrate prior knowledge within an encoder-decoder framework: GFPGAN [36] utilizes a pre-trained face GAN’s rich priors for blind face restoration. Yang et al. [37] integrate a GAN into a U-shaped DNN for enhanced face restoration. Codeformer [36] uses a learned codebook to predict codes, reducing restoration uncertainty. In our approach, we harness a well-trained VQAE as a facial prior to enhance high-frequency restoration from low-quality face images. Through rigorous theoretical analysis and empirical experiments, we establish that the SOVQAE is an effective plug-and-play model for achieving identity-specific, temporally-consistent face restoration, thereby enabling the generation of high-quality talking faces.
2.3 Vector Quantization (VQ) Family
Initially, VQ emerged as a technique within the Signal Processing community [16, 17, 18, 19]. Mathematically, VQ maps signal spaces $\mathbb{R}^{K}$ onto a finite set of Voronoi regions $\{{V_{n}}\}_{n=1}^{N}$, each associated with an anchor vector $\{{\bf c}_{n}\}_{n=1}^{N}\subset\mathbb{R}^{c}$:
$V_{n}=\{{\bf x}\in\mathbb{R}^{c}\mid\|{\bf x}-{\bf c}_{n}\|_{2}\leq\|{\bf x}-{\bf c}_{m}\|_{2},\forall m\neq n\},$ (1)
where $\|\cdot\|_{2}$ denotes the Euclidean distance. Recently, Oord et al. [7] integrated the VQ operation into generative models, sparking a trend of discrete latent representation in the field [20, 15, 21, 22, 23, 24, 25, 13, 26]. For instance, Zhou et al. [20] trained a GAN-based VQAE to learn high-quality (HQ) face priors for blind face restoration. Esser et al. [21] proposed an autoregressive transformer model for high-resolution image generation. MSMC-TTS [22] developed a multi-stage, multi-codebook representation for an advanced text-to-speech system.
Our study regards VQAE as a mechanism for learning and encoding high-frequency facial details, similar to VQ-based approaches. Unlike methods[23, 24, 25, 13, 26] that require additional networks, VQAE inherently possesses noise robustness (see Sec. 3), enabling the recovery of detailed textures from low-quality images. This simplicity makes our method both effective and efficient.
3 Latent Denoising Modeling
In this section, we prove the existence of noise robustness of VQAE via the Lipschitz Continuity theory, where the definition of lipschitz continuity is following:
DEFINITION 3.1.
A function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is called Lipschitz continuous if there exists a constant $L$ such that
$\forall x,y\in\mathbb{R}^{n},~{}\|f(x)-f(y)\|_{2}\leq L\|x-y\|_{2}.$ (2)
The smallest value of $L$ for which the preceding inequality holds is referred to as the Lipschitz constant of $f$. This property guarantees that when a small perturbation $\Delta x$ is applied to $x$, the resulting impact on $f(x)$ is linearly proportional to $\|\Delta x\|$. This characteristic is more stringent than uniform continuity because the Lipschitz constant is independent of specific points in the domain.
As depicted in Fig. 1(b), a VQAE is encapsulated by the quadruple $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$. Here, $\epsilon$ represents the CNN encoder that maps the image domain $\mathbb{R}^{3\times h\times w}$ to the latent space $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The term $\delta$ signifies the CNN decoder, which reconstructs the image from the latent representation in $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The codebook ${\bf C}=\{{\bf c}_{n}\}_{n=1}^{N}\in\mathbb{R}^{c\times N}$ comprises $N$ anchor vectors. Furthermore, $g_{{\bf C}}$ refers to the channel-wise 1-Nearest Neighbor (1-NN) feature matching operation, responsible for the substitution of latent features with their closest anchor vectors within the codebook. As demonstrated in recent studies [9, 8], the following conclusion can be drawn:
THEOREM 3.2.
Given a trained VQAE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, the $\epsilon:~{}\mathbb{R}^{3\times h\times w}\to\mathbb{R}^{c\times h_{o}\times w_{o}}$ is a map with lipschitz continuity such that
$\forall x,y\in\mathbb{R}^{3\times h\times w},~{}\|\epsilon(x)-\epsilon(y)\|_{F}\leq L_{\epsilon}\|x-y\|_{F},$ (3)
where $\|\cdot\|_{F}$ is the Frobenius norm of matrix and $L_{\epsilon}$ is the Lipschitz constant of $\epsilon$.
We provide the proof in Appendix D. Let ${\bf V}_{up}\subset\mathbb{R}^{T\times 3\times h\times w}$ denote the upscaled version of ${\bf V}_{low}$, where ${\bf V}_{low}$ is the output of Wav2Lip. It is evident that each image in ${\bf V}_{up}$ can be viewed as a degraded HQ image. Given that the primary defect in ${\bf V}_{up}$ is blurring, we naturally assume that this degradation can be modeled as Gaussian blur. Therefore, for any ${\bf I}_{up}\in{\bf V}_{up}$ and ${\bf I}_{high}\in{\bf V}_{high}$, we assume that ${\bf I}_{high}+{\bf\mathcal{N}}={\bf I}_{up}$ always holds, where ${\bf\mathcal{N}}=\text{Gaussian\_blurring}({\bf I}_{high})-{\bf I}_{high}$ represents the domain gap between HQ images and their blurred counterparts. Consequently, by Theorem 3.2, we derive the following results:
$\epsilon({\bf I}_{up})\in\{x\in\mathbb{R}^{c\times h_{o}\times w_{o}}:\|x-\epsilon({\bf I}_{high})\|_{F}\leq L_{\epsilon}\|\mathcal{N}\|_{F}\}.$ (4)
This implies that the $\epsilon(\mathbf{I}_{\text{up}})$ resides within the interior of a hypersphere, with the point $\epsilon(\mathbf{I}_{\text{high}})$ as its center and a radius of $L_{\epsilon}\|\mathcal{N}\|_{F}$, in the measure of the Frobenius norm. Furthermore, for each channel of the latent vector, denoted as $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$ and $\epsilon(\mathbf{I}_{\text{high}})_{i,j}\in\mathbb{R}^{c}$, we have that:
$\|\epsilon({\bf I}_{up})_{i,j}-\epsilon({\bf I}_{high})_{i,j}\|_{2}<L_{\epsilon}\|\mathcal{N}\|_{F}.$ (5)
This further indicates an important and evident fact:
THEOREM 3.3.
In a trained VQAE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, we define $d_{C}={\rm min}\{\|{\bf c}_{i}-{\bf c}_{j}\|_{2}~{}:~{}{\bf c}_{i},{\bf c}_{j}\in{\bf C}~{}and~{}i\neq j\}$ as the minimal distance between anchors of codebook ${\bf C}$, $\gamma$ is the maximal distance of training image latent to closest anchor in all channel, and $L_{\epsilon}$ is the Lipschitz constant of $\epsilon$ when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ holds, we have:
$g_{\bf C}(\epsilon({\bf I}_{up}))=g_{\bf C}(\epsilon({\bf I}_{high})).$ (6)
This is what we refer to as the latent denoising mechanism of the VQAE. Specifically, the anchors $\{\mathbf{c}_{n}\}_{n=1}^{N}$ partition the latent domain into $N$ high-dimensional Voronoi regions as defined by Equation (1). Let the nearest anchor to $\epsilon(\mathbf{I}_{\text{high}})_{i,j}$ be $\mathbf{c}_{n}$, and $V_{n}$ be its corresponding Voronoi region.  Theorem 3.2 ensures that $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$ remains within $V_{n}$, provided that $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ is satisfied, thereby enabling the $g_{\mathbf{C}}$ operation to correctly assign $\mathbf{c}_{n}$ to $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$. This condition is valid for each channel of $\epsilon(\mathbf{I}_{\text{up}})$, and thus  Theorem 3.3 applies, allowing the decoder $\delta$ to generate HQ faces without loss. The complete proof of  Theorem 3.3 is detailed in  Appendix E.
Figure 2: Display of the noise robustness performance in Macron subject. Please zoom in for better visualization. From top to bottom, there are Gaussian blurred video, denoised video by SOVQAE, ground truth video, denoised video and Gaussian noised video. For each video, we pick 12 continuous frames clip to demonstrate the temporal-consistency of our method.
In Figure 2, we evaluate the denoising performance on a brief talking sequence from the Macron subject, comprising 12 frames. We assess two common video degradation types: Gaussian blurring and Gaussian noise. For each frame, a $192\times 192$ pixel area is randomly selected and subjected to these noise conditions. The increase in PSNR from 30.26 to 35.78 following Gaussian blurring confirms the temporal validity of Theorem 3.3 and supports our assumptions regarding noise $\mathcal{N}$. Even under more semantically destructive Gaussian noise, our method adeptly recovers detail and maintains temporal consistency, with PSNR rising from 16.08 to 30.36. These observations underscore the robustness of SOVQAE to latent noise and affirm its effectiveness.
The aforementioned theoretical analysis substantiates the existence of a denoising mechanism within the VQAE. In our experiments, we observe that a mere 5 minutes of training data suffices for achieving generalization. Furthermore,  Theorem 3.3 posits that the noise tolerance capability is positively correlated with the ratio $\frac{d_{C}-2\gamma}{2L_{\epsilon}}$. To augment this ratio without compromising the denoising efficacy, we introduce an effective regularization loss, as delineated in Equation (11).
Figure 3: (a) displays the overview of vanilla Wav2Lip[1]. (b) displays the latent denosing process of SOVQAE and the mechanism of regularization loss. Please zoom in for better visualization.
4 LaDTalk
In this section, we introduce the detailed architecture of our LaDTalk. As illustrated in Fig. 3, LaDTalk contains two parts: a vanilla Wav2Lip model synthesis lip-synchronized audio-driven LQ talking faces; an optimized VQAE temporal-consistently lifts LQ faces to HQ faces.
4.1 Vanilla Wav2Lip
As shown in Fig. 3(a), Wav2Lip[1] generates accurate lip-sync by learning from a well-trained SyncNet, where aligns the generated low-half faces and audio in semantic space.
SyncNet processes five-frame face sequences (focusing on the lower half) and the mel-spectrogram of a 0.2-second audio segment, given that the training videos are recorded at 25 FPS. The architecture of SyncNet comprises one audio encoder and one face encoder, both of which are stacks of 2D convolutional neural networks. SyncNet is trained to discern the synchrony between the input audio and video, thereby aiding the lip generator training in a contrastive learning framework.
To accomplish this, training video segments are stochastically sampled to be either in sync or out of sync with audio segments. The training of SyncNet employs a cosine similarity metric coupled with a binary cross-entropy loss, denoted as $\mathcal{L}_{\text{syncnet}}$. Specifically, let $\mathbf{v}\in\mathbb{R}^{D}$ represent the face embedding and $\mathbf{a}\in\mathbb{R}^{D}$ represent the audio segment embedding. The cosine similarity is initially calculated as
${\rm sim}(\mathbf{v},\mathbf{a})=\frac{(\mathbf{v}\cdot\mathbf{a})}{\|\mathbf{v}\|_{2}\|\mathbf{a}\|_{2}},$ (7)
where ()($\cdot$)( ⋅ ) signifies the dot product of vectors. Subsequently, the binary cross-entropy loss is defined as
$\mathcal{L}_{\text{syncnet}}=-y\cdot\log({\rm sim}(\mathbf{v},\mathbf{a}))+(1-y)\cdot\log(1-{\rm sim}(\mathbf{v},\mathbf{a})),$ (8)
where ${\rm sim}(\mathbf{v},\mathbf{a})$ denotes the likelihood that the input audio-video pair is synchronized. A trained SyncNet is then utilized to impose a penalty on the lip generator for any inaccurate generation during the training phase.
The Lip Generator, depicted in Fig. 3(a), comprises three main components: an audio encoder $E_{a}$, a face encoder $E_{f}$, and a face decoder $D_{f}$. The audio encoder $E_{a}$ converts mel-spectrograms of audio into audio features $f_{\text{audio}}\in\mathbb{R}^{D}$. The face encoder $E_{f}$ encodes five reference face frames and five pose face frames into face features $f_{\text{face}}\in\mathbb{R}^{D}$. The reference faces encapsulate essential identity information, such as teeth, lips, and other textures, while the pose faces define the head poses. It is important to note that the lower halves of the pose faces are masked to prevent misleading information regarding the lip region. The unmasked pose faces are the ground truth faces that align with the input audio segment. Additionally, the intermediate features from $E_{f}$ are integrated into the face decoder $D_{f}$ through residual connections [38]. The lowest-level feature of $D_{f}$ is the concatenation of $f_{\text{face}}$ and $f_{\text{audio}}$.
The generator is trained to minimize the L1 reconstruction loss between the generated frames $\mathbf{F}_{\text{gen}}$ and the ground-truth frames $\mathbf{F}_{\text{GT}}$, given by:
$\mathcal{L}_{\text{recon}}=\|\mathbf{F}_{\text{gen}}-\mathbf{F}_{\text{GT}}\|_{1}.$ (9)
Furthermore, to leverage the pretrained SyncNet, five continuous lower half faces and the corresponding 0.2-second audio segment are fed into the pretrained SyncNet. This process minimizes the following expert sync-loss:
$\mathcal{L}_{\text{sync}}=\mathcal{L}_{\text{syncnet}|y=1},$ (10)
where $\mathcal{L}_{\text{sync}}$ ensures that the input audio-video pair is synchronized. This contrastive learning strategy has proven to be effective. According to the statistics from recent papers [2, 14, 3, 39], Wav2Lip continues to achieve the best lip-synchronization performance.
4.2 Space Optimized VQAE
The primary limitations of Wav2Lip are its low resolution ($96\times 96$) and the associated blurring effects. To address these issues, the SOVQAE is designed to consistently upscale LQ faces to HQ faces. Consequently, the SOVQAE should act as a lossless compression technique for talking faces, equipped with a robust noise tolerance capability.
To overcome the resolution limitation, we employ face detection using S3FD [45] to extract facial regions from a talking head video of a specific identity. These extracted regions are subsequently resized and used to train the VQAE. This method allows the codebook to concentrate on capturing the high-frequency facial components. In inference, for given LQ faces, we upscale them to $256\times 256$ before subjecting them to the trained VQAE for latent denoising.
According to Theorem 3.3, perfect latent denoising is achievable when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$. Note that $\gamma$ is much close to zero, because $\gamma$ is optimized via VQ loss (see Eq. 12) in training. Hence, it may seem optimal to maximize $d_{C}$ while minimizing $L_{\epsilon}$. However, this is theoretically infeasible because $L_{\epsilon}$ is positively correlated with the operation norm of convolutional kernels, implying that minimizing $L_{\epsilon}$ equates to L2/L1 parameter regularization[40]. Such regularization would globally compress the latent space distribution, consequently compressing the distribution of anchors in the codebook $\mathbf{C}$ and reducing $d_{C}$ and degrade the noise robustness.
The collect optimization is that increase $d_{C}$ while maintaining $L_{\epsilon}$. In experiments, we find a practical solution is to locally extend the minimum $d_{C}$ with a given lower bound. Thus, the codebook regularization loss we propose is:
$\mathcal{L}_{\text{regular}}=\|d_{C}-\theta\|_{2},$ (11)
where $\theta$ represents the given upper bound. This regularization strategy effectively improves local noise robustness while maintaining global robustness, which will be validated via ablation studies. In addition to this regularization loss, the complete loss function of SOVQAE comprises L2 reconstruction loss, VQ loss[21], perceptual loss, and GAN loss with a patch-based discriminator[58]. The VQ loss is following:
$\mathcal{L}_{\text{VQ}}(\epsilon,\delta,\mathbf{C})$ $=\|x-\hat{x}\|^{2}+\|\text{sg}[\epsilon(x)]-g_{\textbf{C}}(\epsilon(x))\|_{2}^{2}$ (12)
$+\|\text{sg}[g_{\textbf{C}}(\epsilon(x))]-\epsilon(x))\|_{2}^{2},$
where $x$ is the input image, $\hat{x}=\delta(g_{\textbf{C}}(\epsilon(x)))$ is the reconstructed image, $\text{sg}(\cdot)$ denotes the stop-gradient operation following[21]. The later two components of  Eq. 12 are utilized to optimized the $\gamma$.
5 Experiments
Dataset. For comparative analysis, we employ the same set of meticulously edited video sequences utilized in previous works [3, 39, 14], encompassing both English and French dialogues. The average duration of these videos is approximately 8,843 frames. We refer to this collection as the Usual Dataset to differentiate it from our proprietary HF videos. These videos are characterized by their rich textures and high resolution. The average length of these videos is around 11,000 frames. More details of them can be found in supplemental materials. All videos are centered on individual characters and captured at a frame rate of 25 FPS. Consistent with prior research, we adopt a train-to-test split ratio of $10:1$. We empirically set $\theta=1$.
Comparison Baselines. We conduct a comparative analysis of our method with a total of three one-to-all approaches, which include Wav2Lip [1], DINet [12], IP-LAP [51], and three one-to-one NeRF-based methods, namely SyncTalk [2],GeneFace [14] and GeneFace++ [52]. Furthermore, to underscore the significance of our identity-specific face restoration concept, we establish two post-processing baselines using Codeformer [20] and GFPGAN [36].
Evaluation Metrics. To assess the quality of the generated images, we employ the Peak Signal-to-Noise Ratio (PSNR) and the Frechet Inception Distance (FID) [46] as metrics. For evaluating lip synchronization, we utilize the Lip Sync Error Confidence (LSE-C) and Lip Sync Error Distance (LSE-D) scores. Furthermore, to assess the generalizability of our methods, we conduct additional tests using a set of out-of-domain (OOD) audio. This OOD audio comprises 10 diverse voice samples, spanning various languages and genders.
Sliding Evaluation. In our experiments, we observe that the generated video lengths from various methods [1, 12, 14, 2, 39] are not only disparate but also consistently shorter than those of the ground truth videos. Directly assessing image quality in conjunction with video duration could potentially diminish the perceived performance of these methods. To address this issue, we introduce a Sliding Evaluation technique. As illustrated in  Fig. 4, this approach treats the generated video sequence as a dynamic sliding window, calculating the average image quality across all frames captured within the window. The window progressively moves along the timeline, and the maximum value obtained from these iterations is adopted as the final evaluation outcome. We apply this Sliding Evaluation method to all PSNR and FID results to evaluate different methods, thus ensuring a more holistic and equitable assessment.
Figure 4: Illustration of our Sliding Evaluation method.
5.1 Quantitative Comparison
For a more exhaustive evaluation, our quantitative assessment is divided into two segments: 1) Video quality and lip-synchronization comparison, which evaluates image quality frame-by-frame and lip-synchronization performance under original audio-driven conditions, following prior researches [2, 14]; 2) Out-of-Distribution (OOD) audio-driven comparison, where we contrast lip-synchronization with SOTA methods to underscore the superior cross-audio generalizability of our approach, a critical application scenario. The former assesses identity preservation, while the latter evaluates audio generalization capabilities.
Table 1: The quantitative comparison of different methods on Usual Dataset and HF videos. * indicates post-process baselines. Please zoom in for better visualization.
Method Usual Dataset HF videos
[…]
Wav2Lip[1] 34.0979 10.8503 9.077 5.8769 34.6593 17.5281 8.8496 6.8712
DINet[12] 33.9108 9.8385 7.1951 7.4343 34.4849 10.5749 8.3541 7.2340
IP-LAP[51] 34.7263 10.1703 5.1736 8.8160 35.1595 9.9438 8.2604 7.1838
GeneFace[14] 24.8165 21.7084 5.195 - - - - -
GeneFace++[52] 31.1164 20.5506 6.8916 7.5014 33.9692 22.3928 4.6112 11.0401
SyncTalk[2] 36.0574 6.4855 7.054 7.282 39.4050 6.5370 7.816 7.715
CodeFormer*[20] 33.2441 26.3567 9.1739 5.7720 33.7958 16.7215 7.9409 7.2582
GFPGAN*[36] 33.9803 16.7267 9.0417 5.8144 33.7780 17.2715 8.7940 6.5193
Ours 38.9738 3.4505 8.1274 6.44 39.5454 4.3065 8.1392 6.8438
Best; Second best.
Comparison with End-to-End Methods. As detailed in  Tab. 1, on the Usual Dataset, our method secures the highest scores for both PSNR (38.9738) and FID (3.4505), signifying that we achieve state-of-the-art image quality for generated videos. This advancement is credited to the SOVQAE, which retains high-frequency texture details in the codebook and consistently recovers facial textures from low-quality faces. In terms of lip-synchronization, we outperform all methods (excluding our base model), a testament to Wav2Lip’s robust audio-lip alignment capability. Although our method moderately reduces the lip-synchronization performance of the base model, the substantial enhancement in visual quality justifies this trade-off as valuable and insightful. Similar outcomes are observed in the HF videos.
Comparison with Post-Process Methods. For a comprehensive analysis, we also juxtapose our method with recent blind face restoration SOTA techniques. As illustrated in  Tab. 1, our method surpasses two post-process baselines in video quality. This advantage stems from the fact that Codeformer and GFPGAN are tailored for blind face restoration, and their face prior learning from out-of-distribution multi-face datasets results in information leakage, leading to identity errors. Moreover, their image-by-image training strategy is not optimal for video scenarios, causing frame-wise flickering, which significantly detracts from the final video fidelity. Although they maintain the lip-synchronization performance of the base model, their subpar visual fidelity limits their applicability to talking head scenarios. Furthermore, for fair comparison, we also finetune GFPGAN model in ID-specific videos. As shown in Fig. 5, while fine-tuning improves its performance in visual quality, its synchronization decreases. On the other hand, our method has least GPU need (8.5 GPU@4090 hours), best visual quality and comparable synchronization. These findings strongly affirm the efficacy and efficiency of our post-process approach for talking head tasks.
Figure 5: Comparison with fine-tuned GFPGAN model.Note that the sudden collapse of lip synchronization of GFPGAN is caused by the overfitting. We display videos in our supplemental materials.
Table 2: There are 10 cross-lingual and cross-gender 20 second audios in OOD audio-driven experiment. We use different audios drive same subject and calculate LSE-C and LSE-D metrics.
Method DINet[12] IP-LAP[51] GeneFace++[52] SyncTalk[2] Ours
[…]
1 4.7768 8.9890 4.4024 9.3161 4.8974 8.9521 4.2177 9.28129 6.2354 8.0987
2 4.3692 9.0287 4.1531 8.6423 4.6195 8.7908 4.6328 8.0043 5.9480 7.6968
3 5.6343 8.6510 4.8950 9.1588 5.1553 8.8068 6.3866 8.0077 7.4764 7.5445
4 4.2127 8.5808 4.8710 8.6493 3.9984 8.2326 5.7213 7.8372 6.7817 7.5449
5 6.1828 8.5279 3.8994 9.4514 5.9873 8.5823 4.4513 8.8178 6.0246 7.9151
6 6.2382 7.9339 5.1737 8.5354 5.9934 7.7590 6.0401 7.1978 7.1988 7.3960
7 4.6975 8.7730 4.6781 9.2968 5.0774 8.3832 5.4807 8.2488 7.2357 7.1888
8 5.5323 8.3505 4.5502 8.9811 5.9367 7.8090 5.4721 7.9490 6.8894 7.2750
9 5.0672 8.8984 4.7241 9.2790 5.5664 8.6923 5.7529 7.7905 6.9524 7.5649
10 5.5989 8.1889 2.4123 9.2861 6.2121 7.9188 2.9075 7.3401 3.7788 8.4436
average 5.2310 8.5922 4.3759 9.0596 5.3445 8.3927 5.1063 8.0474 6.4521 7.6668
Best.
Comparison on OOD Audio Driving. In the industry, voice generalization is crucial for talking head models. While methods like SyncTalk, GeneFace, and GeneFace++ have made progress, their capabilities still fall short of application requirements. To evaluate our method’s voice generalization, we compared it with recent SOTA models on the OOD audio-driving task. As shown in Table 2, Our method leads in both LSE-C and LSE-D metrics, scoring 6.4521 in LSE-C and 7.6668 in LSE-D, demonstrating strong generalization across diverse voices. It significantly outperforms other methods such as SyncTalk, GeneFace++, IP-LAP, and DINet.
Figure 6: In this qualitative analysis, we compare facial synthesis across various methodologies. The comparative visualization presents a curated selection of discrete keyframes and corresponding sub-frames of lower-half faces, each illustrating the performance of different methods on distinct identities. For an optimized visual assessment, we encourage the reader to zoom in on the figures for enhanced clarity. Please zoom in for better visualization.
Figure 7: Comparison on HF videos. Except for the collapse and lip error (see red boxes), NeRF-based methods also has limitation in stability of face scale. See supplemental materials for more details.
These observations suggest that our method effectively inherits the SOTA audio generalization capability of Wav2Lip and validate the feasibility of the Wav2Lip-based post-process approach. Note, we also conduct experiment with different choices of base models in supplemental materials to demonstrate the plug-and-play performance of LaDTalk framework.
5.2 Qualitative Comparison
To provide a more intuitive assessment of image quality, we present a comparative analysis of our method alongside alternative approaches in  Fig. 6. Specifically, we contrast our method with both end-to-end methodologies, including Wav2Lip [1], SyncTalk [2], IP-LAP [51], and DINet [12], as well as two post-processing techniques, GFPGAN [36] and CodeFormer [20].
In Fig. 6, our method outperforms SyncTalk, a state-of-the-art NeRF-based approach, by better capturing high-frequency facial details. SyncTalk’s struggle to maintain these details, particularly in dynamic scenes, results in blurred mouth and teeth regions. This is attributed to NeRF’s limited expressivity for high-frequency details. In contrast, our method, utilizing SOVQAE, effectively restores individual-specific textures, enhancing fidelity. Compared to Wav2Lip, our method not only preserves lip shape but also achieves more realistic facial textures, indicating superior accuracy. The 7th and 8th columns of the figure, where our method’s identity-specific restoration approach is superior to the two post-processing baselines, which alter the subject’s identity due to information leakage from a large OOD faces dataset [47]. This underscores the necessity of our identity-specific restoration approach. Besides, in more challenging HF videos, our method has more stable performance. In contrast, NeRF-based methods occur geometry collapse (see the 3-th cloumn in Fig. 7) and lip error, when scale their resolution to higher.
In summary, the qualitative comparison underscores that our method surpasses all compared methods in terms of facial texture and lip-synchronization, thereby affirming the efficacy of our post-processing strategy. Note that partly qualitative comparison of HF videos can be found in Appendix C.
5.3 Ablation Study
In this section, we conduct an ablation study to further substantiate the indispensability of each component within our pipeline.
Figure 8: Ablation study of the regularization loss. Without this loss will lead to random noises ( see red boxes ).
VQ Regularization Loss. As previously discussed, in practical applications, training without regularization on the VQ codebook can result in a decrease in noise robustness due to the stochastic optimization of the codebook. To illustrate the efficacy of  Eq. 11 in stabilizing this robustness, we conducted a comparison between the VQAE with and without codebook optimization. As shown in  Fig. 8, the absence of our regularization loss introduces numerous artifacts in the final video, attributable to errors in latent matching.
To further demonstrate the necessity and efficacy of our VQ regularization loss, we conduct detailed ablation studies on the lower bound and optimization settings. As shown in Tab. 3, optimizing the minimal distance pair in the codebook yields the best results: (1) Optimizing pairwise distances within the codebook leads to codebook collapse, as indicated by the first two rows. (2) In contrast, local optimization with $\theta=1$ is effective. (3) Increasing the distance ($\theta=2$) in local optimization degrades performance due to over-constraint. This contrasts with general principles for training VQ-VAEs, where maintaining a sufficiently large distance in the codebook is crucial for diversity in image representation. However, our goal differs: while VQ-VAE aims for diverse codebooks to handle varied images, SOVQAE focuses on enhancing noise robustness for identity-specific faces. We do not explore $\theta<1$ since $d_{c}<1$ is typically the case.
Table 3: Ablation experiments in Usual Dataset. ‘Reg. Obj.’ indicates the regularized object in our regularization loss. ‘Minimal Dist.’ is the $d_{c}$. ‘Average Dist.’ indicates the average distances of each pair of codebook vectors.
[…]
31.21 29.34 4.35 9.58 Average Dist. 2
34.18 18.34 5.67 8.02 Average Dist. 1
34.56 19.27 5.35 8.26 Minimal Dist. 2
38.97 3.45 8.13 6.44 Minimal Dist. 1
Length of training videos. An essential inquiry pertains to the optimal length of video required to train a sufficiently expressive SOVQAE to support our method in achieving realistic and temporally consistent denoising performance. To address this question, we experimentally varied the length of the training video for the SOVQAE and assessed the corresponding impact on performance. The findings, as detailed in  Tab. 4, yield the following insights: Using extended training videos improves the audio-driven capabilities of SOVQAE. A 5-minute training video is adequate for generating realistic talking head videos with our method, indicating similar data requirements to NeRF-based methods. To make our manuscript more sufficient, we introduce our limitations in Appendix B.
Table 4: Ablation study on length of training video. Note that we only compute the PSNR for the Mouth region in this experiment. We conduct this experiment on Macron for more comprehensive study, since it is the longest video in Usual dataset.
Length of video 1 2 3 4 5 6 Full (8’40”)
[…]
6 Conclusion
In this paper, we introduce LaDTalk, a novel method for ID-specific talking head generation that achieves state-of-the-art performance. By leveraging the theory of Lipschitz Continuity, we have theoretically established noise robustness in VQAEs. Building on this foundation, we propose SOVQAE, a plug-and-play denoising model designed to achieve temporally consistent recovery of high-frequency details from low-quality outputs, such as those generated by Wav2Lip. Through extensive experimental validation, we demonstrate that LaDTalk attains new SOTA video quality and out-of-domain lip synchronization accuracy. Additionally, we showcase the versatility of SOVQAE across various base models. This work highlights the potential of post-processing paradigms in the domain of talking head generation. We hope that our contributions will provide valuable insights to the research community.
Supplementary Material
Appendix A Implementation Details.
In our experiments, we configured the codebook with $c=256$ channels and a size of $N=1024$. For the Usual Dataset, we trained the optimized VQAE on face crops resized to a resolution of $256\times 256$ pixels, completing 50 epochs using an NVIDIA 4090 GPU, which amounted to 8.5 GPU hours. For the HFTK Dataset, the face crops were resized to a higher resolution of $512\times 512$ pixels, and the model was trained for 50 epochs on the same hardware, consuming 34 GPU hours. We utilized a downsampling factor of $h_{o}=h/16$ and $w_{o}=w/16$. Additional details regarding the network architecture are presented in our Appendix F. The foundational Wav2Lip model employed in our experiments is a pretrained version sourced from the official repository [^2] https://github.com/Rudrabha/Wav2Lip. For the regularization loss defined in  Eq. 11, we have empirically set $\theta=1$. The high frequency videos consists of 4 4K videos from YouTube, reprocessed to $1024\times 1024$ at 25 fps with a 10M bitrate (compared to 3M in the Usual dataset). They have higher resolution and richer high-frequency textures, as shown in Fig. 9. These videos enable evaluation of identity-specific methods under more detailed and realistic conditions, aligning with industrial needs.
Figure 9: Example display of the processed HF videos. In the left side, we display one keyframe of one of HF videos. In the right side, we display four keyframes of videos in Usual Dataset. It is clear that HF videos have more fine-grained details, sharper textures and larger face region, which are more challenging.
Appendix B Discussion
Limitations. In comparison with existing methods, our approach has a notable limitation in terms of computational requirements for both training and inference phases. While NeRF-based methods such as SyncTalk [2] and GeneFace++ can achieve real-time inference, our method operates at a 1:1.2 inference ratio on an Nvidia 4090 GPU when processing videos at 25 FPS, which indicates that it does not yet meet the criteria for real-time performance. Additionally, our training demands are more substantial; for instance, to train our model on 5 minutes of talking head video data, we require approximately 8.5 GPU hours using an NVIDIA 4090 GPU. These computational demands highlight areas for potential future optimization.
Boarder Impact. In contrast to traditional end-to-end frameworks, our method generates realistic talking head videos through a temporally-consistent post-processing approach. From an alternative viewpoint, we also highlight the potential of harnessing the robust lip synchronization capabilities of a pretrained Wav2Lip model. This suggests that integrating a pretrained Wav2Lip model into a NeRF-based talking head pipeline could significantly enhance per-frame stability and synchronization performance. Moreover, a primary contribution of our work is the theoretical proof and practical demonstration of the noise robustness inherent in the Vector Quantization mechanism. Consequently, it is a logical next step to explore the application of this concept within the context of adversarial attacks [48, 49, 50] to bolster the security and reliability of neural networks.
Future works. To the limitation on computational demand, we consider to explore more efficient network design, which guarantees the noise robustness while decrease the size of whole pipeline, aiming to real-time inference and low training burden. Except for the pipeline optimization, we plan to explore the way of prior integration in talking head task. Specifically, the LQ talking head results of Wav2Lip model are seem as a kind of intermediate representation (like the face keypoints representation in GeneFace and GeneFace++). Following these works’ idea, integrating such more semantic-rich intermediate representation into NeRF or 3DGS pipeline seems like a promising way to achieve better audio generality. Besides, we also notice that SOVQAE incurs decrease of lip-synchronization. Hence, alleviating this phenomena via more delicate network design is also our goal.
Appendix C Additional Experiments
To demonstrate the plug-and-play performance of our framework, we integrated SOVQAE into various base models. As shown in Tab. 5, SOVQAE improves visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Notably, LaDTalk achieves the best overall performance by leveraging the exceptional audio-lip synchronization capabilities of Wav2Lip. This superior synchronization positively impacts PSNR and FID values, thereby enhancing overall video quality and ensuring LaDTalk’s top performance.
Additionally, we provide further visual evidence in Fig. 10. Our SOVQAE significantly enhances high-fidelity details across different ID-specific base models, as evidenced by more detailed forehead wrinkles and finer tooth textures. These improvements highlight the effectiveness of SOVQAE in preserving fine-grained details while maintaining temporal consistency.
Figure 10: Visualization of the Plug-and-Play Experiment. From top to bottom, the figure shows: close-ups of the forehead in the first keyframe from different methods, the first keyframe, the second keyframe, and close-ups of the teeth in the second keyframe.
Table 5: Additional experiments in Usual Dataset.
[…]
GeneFace++ 31.11 20.55 6.89 7.50
GeneFace++ + SOVQAE 33.27 18.21 6.42 8.22
SyncTalk (CVPR 2024) 36.06 6.49 7.05 7.68
SyncTalk + SOVQAE 37.16 5.12 6.58 7.88
Ours 38.97 3.45 8.13 6.44
Appendix D Proof of Theorem 3.1
D.1 Notation
For all positive integer $n$, the set $[n]=\{1,2,\cdots,n\}$.
[…]
For all finite index sets $I$ and $J$, an $I\times J$ matrix $M$ over a ring $R$ is a $|I|\times|J|$ matrix whose rows is indexed by $I$ and whose columns is indexed by $J$, and for all $i\in I,j\in J$, the $(i,j)$-entry of $M_{ij}\in R$.
Let $M$ and $N$ be $I\times J$ and $J\times K$ matrices over $R$ respectively. The product of $MN$ is a $I\times J$ matrix over $R$ whose $(i,k)$-entry is
$(MN)_{ik}=\sum_{j\in J}M_{ij}N_{jk}.$
For all $I\times J$ matrix $M$ over $\mathbb{R}$, the norm $\|M\|$ of $M$ is the Frobenius norm of matrix, that is,
$\|M\|_{F}=\sqrt{\sum_{(i,j)\in I\times J}M^{2}_{ij}}.$
D.2 One layer CNN
We main consider the CNN for graphs, hence we always assume that the shape of input is $c_{i}\times h\times w$.
DEFINITION D.1.
A convolutional layer $\mathcal{L}$ is a data $(\ker\mathcal{L},p\in\mathbb{N}^{2},s\in\mathbb{N}^{2})$ which is consisted of
  1. 1.
    A tensor $\ker\mathcal{L}$ of shape $c_{i}\times c_{o}\times k_{h}\times k_{w}$, whish is called the convolution kernel of $\mathcal{L}$. Where $c_{i}$ is the number of input channels and $c_{o}$ is the number of output channels.
  2. 2.
    A pair $p=(p_{h},p_{w})$ is the padding of $\mathcal{L}$.
  3. 3.
    A pair $s=(s_{h},s_{w})$ is the stride of $\mathcal{L}$.
For convenience, we assume that $s_{h}$ and $s_{w}$ is a factor of $(h-k_{h}+p_{h})$ and $(w-k_{w}+p_{w})$ respectively.
For all convolutional layer $\mathcal{L}=(\ker\mathcal{L},p,s)$, there is a map
$F_{\mathcal{L}}:\mathbb{R}^{c_{i}\times h\times w}\rightarrow\mathbb{R}^{c_{o}\times o_{h}\times o_{w}}$
corresponding to $\mathcal{L}$, where $o_{h}=1+\frac{h-k_{h}+p_{h}}{s_{h}}$ and $o_{w}=1+\frac{w-k_{w}+p_{w}}{s_{w}}$.
The domain $\mathbb{R}^{c_{i}\times h\times w}$ should be considered as $c_{i}$ input channels of shape $h\times w$. Similarly, the codomain $\mathbb{R}^{c_{o}\times h\times w}$ should be considered as $c_{o}$ output channels of shape $o_{h}\times o_{w}$.
Let $V$ and $W$ be $R$-modules. We denote ${\operatorname{Hom}}_{R}(V,W)$ as the space of $R$-linear maps from $V$ to $W$, which is also an $R$-module. And if
$V=\bigoplus_{i\in I}V_{i},\quad\text{and}\quad W=\bigoplus_{j\in J}W_{j}$
then there is an $R$-isomorphism
${\operatorname{Hom}}_{R}\left(\bigoplus_{i\in I}V_{i},\bigoplus_{j\in J}W_{j}\right)\cong\bigoplus_{(i,j)\in I\times J}{\operatorname{Hom}}_{R}(V_{i},W_{j}).$
We refer to [28, Part 3] for details.
The map $F_{\mathcal{L}}$ is linear, then since
$\mathbb{R}^{c_{i}\times h\times w}=\bigoplus_{i=1}^{c_{i}}\mathbb{R}^{h\times w}\quad\text{and}\quad\mathbb{R}^{c_{o}\times o_{h}\times o_{w}}=\bigoplus_{j=1}^{c_{o}}\mathbb{R}^{o_{h}\times o_{w}},$
there is a bijection
${\operatorname{Hom}}_{\mathbb{R}}\left(\mathbb{R}^{c_{i}\times h\times w},\mathbb{R}^{c_{o}\times o_{h}\times o_{w}}\right)=\\ {\operatorname{Hom}}_{\mathbb{R}}\left(\bigoplus_{i=1}^{c_{i}}\mathbb{R}^{h\times w},\bigoplus_{j=1}^{c_{o}}\mathbb{R}^{o_{h}\times o_{w}}\right)\cong\bigoplus_{i,j}{\operatorname{Hom}}_{\mathbb{R}}\left(\mathbb{R}^{h\times w},\mathbb{R}^{o_{h}\times o_{w}}\right)$
the linear map $F_{\mathcal{L}}$ can be identified as $c_{o}\times c_{i}$ many matrix over $\mathbb{R}$ such that the $(i,j)$-matrix are linear maps
$F_{\mathcal{L},i,j}:\mathbb{R}^{h\times w}\rightarrow\mathbb{R}^{o_{h}\times o_{w}}.$
The linear map $F_{\mathcal{L},i,j}$ is the composition of
[…]
where $\iota$ is a linear map represented by a $([h+p_{h}]\times[w+p_{w}])\times([h]\times[w])$ matrix whose $((a,b),(c,d))$-entry is
$\iota_{(a,b),(c,d)}=\begin{cases}1,&\text{if }c=a+p_{h},d=b+p_{w}\\ 0,&\text{otherwise}\end{cases}$
We can check that the map $\iota$ is an isometry embedding.
The linear map $\hat{F}_{\mathcal{L},i,j}$ can also be represented by $([o_{h}]\times[o_{w}])\times([h+p_{h}]\times[w+p_{w}])$ matrix such that the $((a,b),(c,d))$-entry is
$F_{\mathcal{L},i,j,(a,b),(c,d)}=\begin{cases}(\ker\mathcal{L})_{i,j,x,y},&\text{if }(\star)\\ 0,&\text{otherwise}\end{cases}$
where $(\ker\mathcal{L})_{i,j,x,y}$ is the $(i,j,x,y)$-entry in the tensor $\ker\mathcal{L}$, and the condition
$(\star):a=(c-1)k_{h}+x,b=(d-1)k_{w}+y,1\leq x\leq k_{h},1\leq y\leq k_{w}.$
Now we use $\|M\|_{\operatorname{op}}$ to represent the operation norm of the matrix $M$. That is,
$\|M\|_{\operatorname{op}}=\max_{\|x\|=1}\|Mx\|_{2}.$
By definition,
$F_{\mathcal{L}}=\hat{F}_{\mathcal{L}}\circ\iota,$
so we have
$\|F_{\mathcal{L}}\|_{\operatorname{op}}\leq\|\hat{F}_{\mathcal{L}}\|_{\operatorname{op}}\|\iota\|_{\operatorname{op}}=\|\hat{F}_{\mathcal{L}}\|_{\operatorname{op}}.$
Thus, to estimate the operation norm $\|F_{\mathcal{L}}\|$ of $F_{\mathcal{L}}$, it suffices to find the operation norm of $\|\hat{F}_{\mathcal{L}}\|$. Now we first prove a simple but useful lemma.
LEMMA D.2.
Let $\mathcal{A}$ be an $M\times N$ matrix that is be identity with an $m\times n$ block matrix $(A_{ij})_{ij}$, then
$\|\mathcal{A}\|_{\operatorname{op}}\leq\sqrt{mn}\max_{i,j}\|A_{ij}\|_{\operatorname{op}}.$
Proof.
Consider the easier case when $m=1$, for all $x\in\mathbb{R}^{N}$, we split it as $(x_{j})_{1\leq j\leq n}$ which is compatible with the block matrix representation of $\mathcal{A}$. We represent the vector $x_{i}$ by $(x_{jk})_{k}$. Then
$\|x\|_{2}=\sqrt{\sum_{j=1}^{n}\sum_{k}x_{jk}^{2}}=\sqrt{\sum_{j=1}\|x_{j}\|_{2}^{2}}.$
Now $\mathcal{A}x=\sum_{j=1}^{n}A_{1j}x_{j}$, so
$\|\mathcal{A}x\|_{2}=\left\|\sum_{j=1}^{n}A_{1j}x_{j}\right\|_{2}$ $\leq\sum_{j=1}^{n}\|A_{1j}x_{j}\|_{2}$
$\leq\sum_{j=1}^{n}\|A_{1j}\|_{\operatorname{op}}\|x_{j}\|_{2}$
$\leq\sqrt{\sum_{j=1}^{n}\|A_{1j}\|_{\operatorname{op}}^{2}}\sqrt{\sum_{j=1}^{n}\|x_{j}\|_{2}}$
$\leq\sqrt{n}\max_{1j}\|A_{1j}\|_{\operatorname{op}}\|x\|_{2},$
so $\|\mathcal{A}\|_{\operatorname{op}}\leq\sqrt{n}\max_{1j}\|A_{1j}\|_{\operatorname{op}}$.
Then when $n=1$, for all $x\in\mathbb{R}^{N}$, $\mathcal{A}x=(A_{i1}x)_{1\leq i\leq m}$, so by above
$\|\mathcal{A}x\|_{2}=\|(A_{i1}x)^{T}\|_{2}=\sqrt{\sum_{i=1}^{m}\|A_{i1}x\|_{2}^{2}}\leq\sqrt{\sum_{i=1}^{m}\|A_{i1}\|_{\operatorname{op}}^{2}\|x\|_{2}^{2}}\leq\sqrt{m}\max_{i}\|A_{i1}\|_{\operatorname{op}}\|x\|_{2},$
so $\|\mathcal{A}\|_{\operatorname{op}}\leq\sqrt{n}\max_{1j}\|A_{i1}\|_{\operatorname{op}}$.
Now for the general case, denote the $1\times n$ block matrix $\mathcal{A}_{i}=(A_{ij})_{1\leq j\leq n}$, then $\mathcal{A}$ can be identity with the $m\times 1$ block matrix $(\mathcal{A}_{i})_{1\leq i\leq m}$, so by above
$\|\mathcal{A}\|_{\operatorname{op}}\leq\sqrt{m}\max_{i}\|\mathcal{A}_{i}\|_{\operatorname{op}}\leq\sqrt{m}\max_{i}\left(\sqrt{n}\max_{j}\|A_{ij}\|_{\operatorname{op}}\right)=\sqrt{mn}\max_{i,j}\|A_{ij}\|_{\operatorname{op}},$
hence, $\|\mathcal{A}\|_{\operatorname{op}}\leq\sqrt{mn}\max_{i,j}\|A_{ij}\|_{\operatorname{op}}$. ∎
By the above Lemma, we have
$\|\hat{F}_{\mathcal{L}}\|_{\operatorname{op}}\leq\sqrt{c_{i}c_{o}}\max_{i,j}\|\hat{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}.$
Now we will estimate the operation norm of $\hat{F}_{\mathcal{L},i,j}$.
By the definition of $\hat{F}_{\mathcal{L},i,j}$, we have a key observation: we can rearrange the index such that the rows of $\hat{F}_{\mathcal{L},i,j}$ are very similar, the matrix after rearranging is denoted by $\tilde{F}_{\mathcal{L},i,j}$. Moreover, we have
$\|\hat{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}\leq\|\tilde{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}.$
For every vector $v=(v_{i})_{i}\in\mathbb{R}^{n}$ and for all positive integer $m$, we define the shifted vector
$v[m]=(\underbrace{0,0,\cdots,0}_{m\text{ times}},v_{1},v_{2},\cdots,v_{n-m}).$
After a proper rearrangement of index, let $\wp_{k}$ be the $k$-th row of $\tilde{F}_{\mathcal{L},i,j}$, by calculation,
$\wp_{k}=\wp_{1}[s_{h}(k-1)],$
hence, we denote $\wp_{1}$ by Weierstrass-$\wp$. By the fact
$\|A\|_{\operatorname{op}}=\sqrt{\rho(AA^{T})},$
where $\rho(A)$ is the spectral radius of $A$. See details for [29, Theorem 6.15].
Notice that $\tilde{F}_{\mathcal{L},i,j}\tilde{F}_{\mathcal{L},i,j}^{T}$ is a positive definite symmetric Toeplitz matrix
$\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n}\\ c_{1}&c_{0}&c_{1}&\cdots&c_{n-1}\\ c_{2}&c_{1}&c_{0}&\cdots&c_{n-2}\\ \vdots&\vdots&\vdots&\cdots&\vdots\\ c_{n}&c_{n-1}&c_{n-2}&\cdots&c_{0}\end{pmatrix}$
where $c_{k}=\wp_{1}\cdot\wp_{k+1},n=o_{h}o_{w}-1$.
PROPOSITION D.3.
If $s_{h}\geq k_{h}$ and $s_{w}\geq k_{w}$, then $\|\hat{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}\leq\|\wp\|_{2}=\|\ker\mathcal{L}_{ij}\|_{F}$.
Proof.
By calculation, if $s_{h}\geq k_{h}$ and $s_{w}\geq k_{w}$, then $\wp_{u}\cdot\wp_{v}=0$ for all $u\neq v$, hence $AA^{T}={\operatorname{diag}}(\|\wp_{1}\|^{2},\|\wp_{2}\|^{2},\cdots)$, and by definition of $F_{\mathcal{L}}$,
$\|\wp_{k}\|_{2}=\|\wp\|_{2}=\|\ker\mathcal{L}_{ij}\|_{F},$
hence,
$\|\hat{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}\leq\|\tilde{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}=\sqrt{\rho\left(\tilde{F}_{\mathcal{L},i,j}\tilde{F}_{\mathcal{L},i,j}^{T}\right)}=\sqrt{\|\wp\|_{2}^{2}}=\|\wp\|_{2}=\|\ker\mathcal{L}_{ij}\|_{F},$
we are done. ∎
In practical applications, often $k_{h}\ll h,k_{w}\ll w$, which means that $AA^{T}$ is a positive definite symmetric banded Toeplitz matrix, there are many theorems estimating the operation norm of such matrices. Now we state some relative results.
Let $f$ be a real value function. The essential supremum $M_{f}$ of $f$ is the smallest number such that $f(x)\leq M_{f}$ for all $x$ except on a set of measure 0.
For all Toeplitz matrix
$C=\begin{pmatrix}c_{0}&c_{-1}&c_{-2}&\cdots&c_{-n}\\ c_{1}&c_{0}&c_{-1}&\cdots&c_{-(n-1)}\\ c_{2}&c_{1}&c_{0}&\cdots&c_{-(n-2)}\\ \vdots&\vdots&\vdots&\cdots&\vdots\\ c_{n}&c_{n-1}&c_{n-2}&\cdots&c_{0}\end{pmatrix}$
the Fourier series $f_{C}(\lambda)$ associate to $C$ is defined by
$f_{C}(\lambda)=\sum_{k=-n}^{n}c_{k}\mathrm{e}^{\mathrm{i}k\lambda},$
then we have
$\rho(C)\leq 2M_{|f_{C}|}.$
In particular, if $C$ is Hermitian, that is, $c_{-k}=c_{k}^{*}$, then
$\rho(C)\leq M_{|f_{C}|}.$
We refer to [27, Lemma 4.1] for details.
We denote the Fourier series $f_{\tilde{F}_{\mathcal{L},i,j}\tilde{F}_{\mathcal{L},i,j}^{T}}$ associate to $\tilde{F}_{\mathcal{L},i,j}\tilde{F}_{\mathcal{L},i,j}^{T}$ by $f_{\mathcal{L},i,j}$. The above result gives that
$\|\hat{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}\leq\|\tilde{F}_{\mathcal{L},i,j}\|_{\operatorname{op}}=\sqrt{\rho\left(\tilde{F}_{\mathcal{L},i,j}\tilde{F}_{\mathcal{L},i,j}^{T}\right)}\leq\sqrt{M_{\left|f_{\mathcal{L},i,j}\right|}}.$
Therefore, we concliude that
$\|F_{\mathcal{L}}\|_{\operatorname{op}}\leq\max_{i,j}\sqrt{c_{i}c_{o}M_{\left|f_{\mathcal{L},i,j}\right|}}.$
As a conclusion, if $\mathcal{L}$ is a convolutional layer, and $F_{\mathcal{L}}$ is the map correspoeing to $\mathcal{L}$, then for all $x,y\in\mathbb{R}^{c_{i}\times h\times w}$, we have
$\|F_{\mathcal{L}}(x)-F_{\mathcal{L}}(y)\|_{F}\leq M_{\mathcal{L}}\|x-y\|_{F}.$
where $M_{\mathcal{L}}=\max_{i,j}\sqrt{c_{i}c_{o}M_{\left|f_{\mathcal{L},i,j}\right|}}$.
We always assume that the activation function is Lipschitz continuous and its Lipschitz constant is known.
D.3 Multiple layers CNN
The above results show that single-layer CNN is Lipschitzian continuous with computable Lipschitz constant. Since any CNN network is a composite of multiple single-layer CNN networks and activation functions, we conclude that any CNN is Lipschitzian continuous with computable Lipschitz constant.
We, therefore, reach the following conclusion.
Let $\epsilon:\mathbb{R}^{c_{i}\times h\times w}\rightarrow\mathbb{R}^{c_{o}\times o_{h}\times o_{w}}$ is a multi-layer CNN, $\mathcal{L}_{k}$ the $k$-th convolutional layer, and let the Lipschitz constant of the $k$-th activation functions be $L_{k}$. Then
$\forall x,y\in\mathbb{R}^{c_{i}\times h\times w},\quad\|\epsilon(x)-\epsilon(y)\|_{F}\leq L_{\epsilon}\|x-y\|_{F},$
where $L_{\epsilon}=\prod_{k}M_{\mathcal{L}_{k}}L_{k}$.
Appendix E Proof of Theorem 3.2
Now we mainly consider a CNN as a Lipschitzian continuous function $\epsilon$ with Lipschitzian constant $L_{\epsilon}$. The spaces $\mathcal{S}$ of graphs of size $h\times w$ is a subset of the space $\mathbb{R}^{c_{i}\times h\times w}$. The range $\epsilon(\mathcal{S})$ can be seen as an encoding of $\mathcal{S}$. In this paper, we utilize VQAE to learn suitable encoding of space $\mathcal{S}$ and simultaneously select anchors in this encoding. Our goal is for this network to learn the features of high-quality images and minimize the distance between codebook and the features. Meanwhile, we also wish the network to be robust with mild input disturbance .Next, we will demonstrate the feasibility of our approach.
We first prove it in the case of single channel:
Let $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ be VQAE with single channel latent space, where $\epsilon$ is the CNN encoder with Lipschitzian constant $L_{\epsilon}$, and $\mathbf{C}\subset\epsilon(\mathcal{S})$ is a set of codebook anchors. Define $d_{\mathbf{C}}=\min\{\|a-b\|_{F}:a,b\in\mathbf{C},a\neq b\}$ and $\gamma$ be the maximal distance of high-quality image latent to closest anchor, that is, $\gamma=\max_{p\in\epsilon(\mathbf{HQ})}d(p,\mathbf{C})$, where $\mathbf{HQ}$ is the space of high quality images, and $p\in\mathbb{R}^{c}$ is single channel vector. Assume that $2\gamma<d_{\mathbf{C}}$ (for almost all cases, it is naturally satisfied). For all low-quality image $y$ corresponding to a high-quality image $x$ such that
$\|x-y\|_{F}<\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}},$
then we have
$\|\epsilon(x)-\epsilon(y)\|_{F}\leq L_{\epsilon}\|x-y\|_{F}<\frac{d_{\mathbf{C}}-2\gamma}{2}.$
Let $g_{\mathbf{C}}(\epsilon(x))=s$, that is, $s\in\textbf{C}$ is the anchor closest to $\epsilon(x)$. Then
$\|\epsilon(x)-s\|=d(\epsilon(x),\mathbf{C})\leq\max_{p\in\epsilon(\mathbf{HQ})}d(p,\mathbf{C})=\gamma.$
For anchor $s$,
$\|\epsilon(y)-s\|_{F}$ $=\|\epsilon(y)-\epsilon(x)+\epsilon(x)-s\|_{F}$
$\leq\|\epsilon(y)-\epsilon(x)\|_{F}+\|\epsilon(x)-s\|_{F}$
$<\frac{d_{\mathbf{C}}-2\gamma}{2}+\gamma$
$=\frac{d_{\mathbf{C}}}{2}$
so $\|\epsilon(y)-s\|_{F}<\frac{d_{\mathbf{C}}}{2}$.
For each anchor $s\neq a\in\mathbf{C}$, we claim that $\|\epsilon(y)-a\|\geq\frac{d_{\mathbf{C}}}{2}$: if not, then
$d_{\mathbf{C}}$ $=\min\{\|a-b\|_{F}:a,b\in\mathbf{C},a\neq b\}$
$\leq\|a-s\|_{F}$
$=\|a-\epsilon(x)+\epsilon(x)-s\|_{F}$
$\leq\|\epsilon(x)-a\|_{F}+\|\epsilon(x)-s\|_{F}$
$<\frac{d_{\mathbf{C}}}{2}+\frac{d_{\mathbf{C}}}{2}$
$=d_{\mathbf{C}}$
so $d_{\mathbf{C}}<d_{\mathbf{C}}$, a contradiction. Thus, $s$ is also the anchor closest to $\epsilon(y)$, which means that
$g_{\mathbf{C}}(\epsilon(y))=s=g_{\mathbf{C}}(\epsilon(x)).$
When latent space is multi-channel ($\epsilon(x)\subset\mathbb{R}^{h\times w\times c}$), for all low-quality image $y$ corresponding to a high-quality image $x$ such that
$\|x-y\|_{F}<\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}}.$
Let $\epsilon(x)_{ij},\epsilon(y)_{ij}\in\mathbb{R}^{c}$ be two single-channel vector of $\epsilon(x),\epsilon(y)$, and $s_{ij}=g_{\mathbf{C}}(\epsilon(x)_{ij})\in\textbf{C}$ be the closest anchor of $\epsilon(x)_{ij}$. Then
$\|\epsilon(x)_{ij}-\epsilon(y)_{ij}\|_{F}\leq\|\epsilon(x)-\epsilon(y)\|_{F}<\frac{d_{\mathbf{C}}-2\gamma}{2},$
by the proof of the single channel case,
$g_{\mathbf{C}}(\epsilon(y)_{ij})=g_{\mathbf{C}}(\epsilon(x)_{ij}),$
which shows that
$g_{\mathbf{C}}(\epsilon(y))=(g_{\mathbf{C}}(\epsilon(y)_{ij}))_{ij}=(g_{\mathbf{C}}(\epsilon(x)_{ij})_{ij}=g_{\mathbf{C}}(\epsilon(x)).$
Therefore, we could get the desired high-quality image $x$ corresponding to $y$ after decoding by $\delta$.
REMARK E.1.
If the VQAE $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ is convergent, the distance of high-quality images to codebook $\mathbf{C}$ is extremely small, which means that $\gamma\ll 1$ is negligible.
As a corollary, if $\mathbf{I}_{high}$ is a high-quality image, and $\mathbf{I}_{up}=\mathbf{I}_{high}+\mathcal{N}$ where $\mathcal{N}$ is the Gaussian degradation with $\|\mathcal{N}\|_{F}<\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}}$, then the above shows that
$g_{\mathbf{C}}\left(\epsilon(\mathbf{I}_{up})\right)=g_{\mathbf{C}}\left(\epsilon(\mathbf{I}_{high})\right).$
Therefore, for CNN $\epsilon$, it can correctly match images with a distance of no more than $\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}}$ from the high-quality images. The correctness of this result requires us to ensure that the selected anchors are all high-quality images, which requires us to use high-quality images for training so that the model can extract features from high-definition images. The judgment range of this model is determined by three parts: first, the Lipschitz constant $L_{\epsilon}$ of $\epsilon$, and second the distance $c$ between anchors in $\epsilon(\mathcal{S})$, and third, the distribution of anchors in the original space $\mathcal{S}$. These three factors are interdependent. To be precise, while ensuring that the anchor points taken are all high-quality images, the effective range of the model is
$\bigcup_{a\in\mathbf{C}}B\left(\epsilon^{-1}(a),\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}}\right),$
where $B(a,r)=\{x:\|x-a|<r\}$.
This requires us to ensure that the anchors are all high-quality images, and to make the anchors as uniform as possible in the original space $\mathcal{S}$ if the number of anchors are fixed, while also making the ratio $\frac{d_{\mathbf{C}}-2\gamma}{2L_{\epsilon}}$ as large as possible. It should be noted that the values of $d_{\mathbf{C}}$ and $L_{\epsilon}$ are correlated. For example, we can always multiply by a constant $\alpha<1$ to change the Lipschitz constant to $\alpha L_{\epsilon}$, but at the same time, the parameter $d_{\mathbf{C}}$ and $\gamma$ also becomes $\alpha d_{\mathbf{C}}$ and $\alpha\gamma$ respectively, which implies that the ratio remains unchanged.
In theory, the selection of anchors is independent of CNN $\epsilon$. However, it’s worth noting that as $\epsilon$ changes, the model’s ability to capture features of high-quality images also changes, affecting the selection of anchor points. Having a large distance $d_{\mathbf{C}}$ between anchor points is not ideal, as it affects both the Lipschitz constant of the CNN, as mentioned above, and the distribution of anchors in the space $\mathcal{S}$.
Similarly, we observe that having a Lipschitz constant $L_{\epsilon}$ that is too small is also not ideal, as it may cause the distances between images of different content to be too small, leading to decreased robustness of the model. An ideal scenario is where the distances between images of the same content are compressed while the distances between images of different content are relatively large. Therefore, we choose to impose reasonable requirements on the distribution of anchor points to train and improve the effectiveness of our model.
Due to the characteristics of images, in practical applications, we often process each part of the image locally. Therefore, we apply the above method to each part to achieve more accurate results, and each part can be regarded as a whole. Therefore, our above argument still holds in this case.
Appendix F More implementation details
Each network component is displayed in Table 6. In experiment, we have 4 downsample blocks. hence we have $m=4,f=16$ in table.
Encoder Decoder
$x\in\mathbb{R}^{H\times W\times C}$ $z_{q}\in\mathbb{R}^{h\times w\times n_{z}}$
Conv2D $\to\mathbb{R}^{H\times W\times C^{\prime}}$ Conv2D $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$
$m\times\{\text{Residual Block, Downsample Block}\}\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$ Residual Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$
Residual Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$ Non-Local Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$
Non-Local Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$ Residual Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$
Residual Block $\to\mathbb{R}^{h\times w\times C^{\prime\prime}}$ $m\times\{\text{Residual Block, Upsample Block}\}\to\mathbb{R}^{H\times W\times C^{\prime}}$
GroupNorm, Swish, Conv2D $\to\mathbb{R}^{h\times w\times n_{z}}$ GroupNorm, Swish, Conv2D $\to\mathbb{R}^{H\times W\times C}$
Table 6: High-level architecture of the encoder and decoder of our SOVQAE. The design of the networks follows the architecture presented in [58] with no skip-connections. For the discriminator, we use a patch-based model. Note that $h=\frac{H}{2^{m}},w=\frac{W}{2^{m}}$ and $f=2^{m}$.